Please help me solve the following expression:
Divide:
((x+2)/(x-2))/((x^2-4)/(2-x))
: Please help me solve the following expression:
Divide:
((x+2)/(x-2))/((x^2-4)/(2-x))

Answer by Fombitz(1756)

((x+2)/(x-2))/((x^2-4)/(2-x))=((x+2)/(x-2))/((x^2-4)/(-(x-2)))Take out (-1) from bottom expression denominator.
((x+2)/(x-2))/((x^2-4)/(2-x))=((x+2))/((x^2-4)/(-1))Denominators in top and bottom expression cancel out.
Factor the quadratic and substitute,
x^2-4=(x+2)(x-2)
((x+2)/(x-2))/((x^2-4)/(2-x))=(-(x+2))/((x+2)(x-2))
((x+2)/(x-2))/((x^2-4)/(2-x))=(-(1))/((x-2))
((x+2)/(x-2))/((x^2-4)/(2-x))=1/(2-x)